OpenStudy (anonymous):

A scientist counts 35 bacteria in a culture and finds that the number of bacteria triples each hour. The function y=35*3^x models the number of bacteria after x hours. Graph the function. Use the graph to estimate when there will be about 550 bacteria in the culture.

OpenStudy (phi):

\[ y= 35\cdot 3^x \] pick numbers for x and figure out what y is y is the number of bacteria after x hours. It makes sense x would be zero or bigger. so start with x=0 3^0 is 1 so y= 35*1= 35 so (0,35) is one point can you find more points?

OpenStudy (anonymous):

could you help

OpenStudy (phi):

can you find more points?

OpenStudy (anonymous):

how are you doing this

OpenStudy (anonymous):

triples each hour. The function y = 35 • 3^x models the number of bacteria after x hours. Graph the function. (You do not need to submit the graph). Use the graph to estimate when there will be about 550 bacteria in the culture. There will be 550 in bacteria after 5.2 hours Comment: Try making a table. One x-y pair would be (1, 105). If one hour goes by (x), then there are 105 (y) bacteria in the culture. When will there be about 550 bacteria in the culture? y= 35 . 3^x (35 is the initial number of bacteria) y= 35 . [ 3^(1) (substitute when x=1, so that u can graph it) y=105 when x=1 (these r the points which u r goin to plot in the graph) then y=35.3^x (substite when x=2) y = 35 * 3^2 y = 35 * 9 = 315 then do the same thing and u get a graph of x/y axis, plot the points and extend the curve and then draw a line(dotted) from the y=550 and chck x axis= which is the hours.... ] Expert answered|wiwit|Points 465|

OpenStudy (phi):

3^1 means 3 3^2 means 3*3 3^3 means 3*3*3 to find 3^x pick an x, for example 1, 2, 3 find 3^x then multiply by 35

OpenStudy (anonymous):

ok thanks phi

OpenStudy (anonymous):

ur answer is point 465

OpenStudy (anonymous):

thats what i got thanks abayomi12

OpenStudy (anonymous):

ok